Opened 5 years ago

Closed 3 years ago

#508 closed flight processing (fixed)

GB13/09, flight day 308b/2013, Pennines

Reported by: knpa Owned by: tipo
Priority: alpha 4 medium Milestone: 2013 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by tipo)

Data location: /users/rsg/arsf/arsf_data/2013/flight_data/uk/GB13_09-2013_308b_Pennines/

Data arrived from ARSF via network transfer on 05/11/2013

GB13/09 Scientific objective: Evaluating the influence of error on change detection assessments in heavily disturbed landscapes using repeat laser scanning

GB13/09 Priority: 8

PI: Danny Donoghue

Sensors:

Camera (Delivered 29/12/2013)
Eagle (not requested)
Hawk (not requested)
LiDAR (Delivered 29/12/2013)
FW LiDAR (Delivered 29/12/2013)

Change History (21)

comment:1 Changed 5 years ago by knpa

  • Description modified (diff)

comment:2 Changed 5 years ago by knpa

  • Description modified (diff)

comment:3 Changed 5 years ago by knpa

  • Description modified (diff)

comment:4 Changed 5 years ago by tipo

Found pitch and roll offsets for LiDAR and SCTs for hyperspectral, Ready to reprocess once the boresights are done.

SCTs:

Flightline Eagle Hawk
1 0 -0.05
2 1.9 (no nav) -0.05
3 0 -0.05
4 -0.01 -0.05
5 0.3 -0.05
6 0.99 -0.05
7 0 -0.05
8 0 -0.05
9 0 -0.05
10 0 -0.05
11 0 -0.05
12 0 -0.05
13 0 -0.05
14 0 -0.05
15 2.5 (no nav) -0.05
16 0 -0.05
17 0 2.1 (no nav)
18 0 -0.05
19 0 -0.05
20 0 -0.05
21 1.35 (no nav) -0.05

comment:5 Changed 5 years ago by stgo

Completed RCD processing, awaiting delivery check.

comment:6 Changed 5 years ago by tipo

The higher altitude LiDAR lines (1-6+21) were processed using pitch: -0.003112592 and roll: -0.001805724
The lower altitude lines (7-20) were processed using pitch: -0.003312592 and roll: -0.001455724

comment:7 Changed 5 years ago by tipo

The lower altitude lines appear to have had range issues. They go outside the maximum height as they pass over a valley. The data seems to be ok on most of the lines, except line 10 (131104_153951) which has a large hole where the valley should be.

Last edited 5 years ago by tipo (previous) (diff)

comment:8 Changed 5 years ago by tipo

Checked RCD. A lot of the images have a blue tint. Might be worth mentioning in the data quality section of the readme.

comment:9 Changed 5 years ago by stgo

Have added note to the RCD readme.

comment:10 Changed 5 years ago by stgo

Only 7 of the waveform files have waveform data present in them. I have processed one of the lines which lacked waveform data using the same reg file as tipo, which resulted in success. I am processing the rest of the files on my machine to check the rest.

comment:11 Changed 5 years ago by stgo

Beginning Hyperspectral D/C.

comment:12 Changed 5 years ago by stgo

Hyperspectral D/C finished, looks good.

Add a comment to the Readme about high levels of underflow in hawk lines 1, 5 (in the higher bands), 6, 9, 10, 13, 14, 15, 17, 18, 19 and 21. Otherwise ready.

comment:13 Changed 5 years ago by tipo

Recreated LiDAR delivery with the reprocessed waveform data.

comment:14 Changed 5 years ago by stgo

D/Cd new LiDAR delivery, ready to deliver.

comment:15 Changed 5 years ago by tipo

Delivered RCD and LiDAR 29/11/13

comment:16 Changed 5 years ago by tipo

  • Component changed from Processing: general to Archiving
  • Description modified (diff)

comment:17 Changed 5 years ago by tipo

Added comment to hyperspectral readme.

comment:18 Changed 4 years ago by stgo

Archiving
Starting archiving

comment:19 Changed 4 years ago by stgo

Archiving
Rsyncing to NEODC

comment:20 Changed 4 years ago by stgo

Archiving
Upload to NEODC complete

comment:21 Changed 3 years ago by dac

  • Resolution set to fixed
  • Status changed from new to closed

Archiving

Data now available from NEODC.

Note: See TracTickets for help on using tickets.